OpenStudy (anonymous):

Which is the set of roots for the equation? x2 + 3x - 18 = 0 {18,-1} {-6,3} {-2,9} {9}

OpenStudy (anonymous):

Factor it or use the quadratic formula.

OpenStudy (anonymous):

Finishd using open study. You don got on my last fraking nerve fool.

OpenStudy (anonymous):

If you don't understand, ask. This isn't a site to get answers for your mediocre questions. These can easily be solved, if you're refusing to learn, I suggest you get off this site.

OpenStudy (anonymous):

{-6,3} if you factor it: x2+3x-18 = (x+6)(x-3) =0

OpenStudy (anonymous):

@myko, this moron isn't interested in how to do it, he just wants answers.

OpenStudy (anonymous):

@petewe, kk